# Evaluating Entity Retrieval in Electronic Health Records: a Semantic Gap Perspective

### v1

@@ 5.2. Benchmark Results
The performances of various baseline retrieval methods are shown in Table 2. We observe that the performance rankings of the models are inconsistent across the two retrieval settings. For example, while bge model presented much higher performances over BM25 in Single-Patient Retrieval, its performance in Multi-Patient Retrieval is comparable to, or even lower than BM25. This discrepancy suggests that these two scenarios emphasize different kinds of abilities, highlighting the need for a comprehensive evaluation on both settings.
@@ 5.4. Query type assessment
We also conduct detailed analysis regarding different query types: disease, procedure, and drugs. We consider the same baseline methods as in Section 5.3, and the results are shown in Table 4. In Simple-Patient Retrieval, generally speaking, model performances on disease queries and procedure queries are comparable, but significantly lower than drug queries across all methods. This may be attributed to the unique characteristics of drug queries: most of them are single-word and annotated through exact string match. These features also account for the fact that BM25 ranks top 1 among all retrievers including RRF in terms of shallow metrics in Multi-Patient Retrieval. Specifically, query expansion, despite a slightly higher recall, causes a significant drop in MRR and NDCG@10. Dense retrievers behave suboptimally in drug retrieval under Multi-Patient setting, indicating that using dense representations to match detailed information such as drug mentions may still be a huge challenge. RRF, a combination of sparse and dense retrieval results leads to a decreased shallow metrics and a significant higher recall.
@@ 6. FUTURE RESEARCH DIRECTIONS
Through a comprehensive analysis of the performances of various retrieval methods on the task of EHR retrieval, we point out several potential future research directions. Firstly, we observe that general domain retrievers outperform medical domain ones, highlighting the significance of the discrepancy between EHRs and conventional biomedical training corpus. The advantage of general domain retrievers over BM25 on CliniQ is also moderate, calling for researches tailored for the task of EHR retrieval.

### v2

@@ 5.2. Benchmark Results
The performance of various baseline retrieval methods are shown in Table 3. We observe that the performance rankings of the models are inconsistent across the two retrieval settings. For example, while bge model presented much higher performance over BM25 in Single-Patient Retrieval, its performance in Multi-Patient Retrieval is comparable to, or even lower than BM25. This discrepancy suggests that these two scenarios emphasize different kinds of abilities, highlighting the need for a comprehensive evaluation on both settings.
@@ 5.4. Query type assessment
We also conduct detailed analysis regarding different query types: disease, procedure, and drugs. We consider the same baseline methods as in Section 5.3, and the results are shown in Table 5. In Simple-Patient Retrieval, generally speaking, model performance on disease queries and procedure queries are comparable, but significantly lower than drug queries across all methods. This may be attributed to the unique characteristics of drug queries: most of them are single-word and annotated through exact string match. These features also account for the fact that BM25 ranks top 1 among all retrievers including RRF in terms of shallow metrics in Multi-Patient Retrieval of drugs. Specifically, query expansion, despite a slightly higher recall, causes a significant drop in MRR and NDCG@10. Dense retrievers behave suboptimally in drug retrieval under Multi-Patient setting, indicating that using dense representations to match detailed information such as drug mentions may still be a huge challenge. RRF, a combination of sparse and dense retrieval results leads to a decreased shallow metrics and a significant higher recall.
@@ 6. FUTURE RESEARCH DIRECTIONS
Through a comprehensive analysis of the performance of various retrieval methods on the task of EHR retrieval, we point out several potential future research directions. Firstly, we observe that general domain retrievers outperform medical domain ones, highlighting the significance of the discrepancy between EHRs and conventional biomedical training corpus. The advantage of general domain retrievers over BM25 on CliniQ is also moderate, calling for researches tailored for the task of EHR retrieval.
@@ 7. CONCLUSION
Limitations
This study has several limitations that warrant discussion. First, our benchmark currently focuses exclusively on discharge summaries, while inclusion of diverse clinical note types would better reflect real-world challenges. However, comprehensive annotation of all clinical notes, which vastly outnumber discharge summaries, would be prohibitively expensive for us, with even 1,000 patients potentially requiring tens of thousands of dollars in annotation costs.
Second, although our benchmark covers most commonly searched entities, it omits certain categories such as symptoms and anatomical structures. This limitation stems from the absence of standardized coding systems (comparable to ICD) for these entities, which form the basis of our query collection. Furthermore, without a unified vocabulary system, we cannot effectively cluster identical queries across different patients or note chunks to enhance benchmark construction.
Third, as noted in Section 4.1, the incompleteness of ICD labels introduces unavoidable false negatives into our benchmark. While exhaustive annotation of all 1,246 queries against 16,550 chunks in CliniQ would be computationally intractable (a challenge shared by all retrieval benchmarks, actually), established benchmarks have nonetheless proven valuable for fair evaluation and have significantly advanced retrieval research over time.